OpenStudy (anonymous):

Write an equation for a line perpendicular to 3y-15x=12 and passing through the point (-10,-2)

OpenStudy (anonymous):

rearrange the formula to make y the subject. This allows you to find the slope of the line: \[y=mx+\] where m is the slope. Then since the line is perpendicular use:\[m_1*m_2=-1\] to find the slope of the line you need. Then substitue everything into the equation: \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

no, try again

OpenStudy (anonymous):

Yes, sorry if I didn't see the '.'

OpenStudy (kyosuke):

solution in the image Luck.!

Directrix (directrix):

Once finished with cranking out the Algebra, the online calculator can be used as a checking device. http://calculator.tutorvista.com/perpendicular-line-calculator.html